Find the equation of the line through the points (4, -1) and (2, -7). write in standard form using integers only.
debsyzach@yahoo.com

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

Given 2 points example.
Find the equation of the line thru the points (2,1) and (3,5)
---------------------
This is a 2 step process. First find the slope of the line thru the points.
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
---------
Now use y = mx + b with either point to find b, the y-intercept.
y = mx + b
5 = 4*3 + b
b = -7
-------
y = mx + b
y = 4x - 7 is the answer.
------------------------
Another approach using determinants:
|x y 1|
|2 1 1| = 0
|3 5 1|
--------
x(1-5) - y(2-3) + 1(10-3) = 0
-4x + y + 7 = 0
y = 4x - 7
---------
For further assistance, or to check your work, email me via the thank you note
or at Moral Loophole@aol.com

Answer by solver91311(24713)



Use the two-point form of the equation of a line.



where and are the coordinates of the given points.

Then manipulate the result so that it is in Standard Form, namely



Remember,


John